# Wake-Informed 3D Path Planning for Autonomous Underwater Vehicles Using A* and Neural Network Approximations

### v1

@@ 5.2 Model Architecture
Both neural network architectures are fully connected feedforward networks, designed to regress the sequence of waypoint coordinates directly from the input features. The architecture consists of an input layer, two hidden layers and an output layer as detailed below:
  • Input Layer: Accepts an 8-dimensional input vector comprising the normalized starting position, goal position, flow speed, and flow angle.
  • $1^{st}$ Hidden Layer: 128 neurons with ReLU activation.
  • $2^{nd}$ Hidden Layer: 256 neurons with ReLU activation.
  • Output Layer: Outputs a 390-dimensional vector representing the flattened sequence of waypoint coordinates $(x_{i},y_{i},z_{i})$ for $i=1,\dots,130$.
The choice of layer sizes was determined empirically to balance model capacity and computational efficiency. ReLU activation functions were used to introduce non-linearity and mitigate the vanishing gradient problem [30]. A fully connected architecture was chosen for its simplicity and to serve as a baseline, ensuring that future studies can easily compare advanced architectures (e.g., RNNs, Transformers) against this initial model.
@@ 5.3 Training Procedure
where $x$ is the feature value, $\mu$ is the mean, and $\sigma$ is the standard deviation computed from the training set. The dataset was then split into training and validation sets using an 80/20 split, resulting in 14,400 training samples and 3,600 validation samples. The split was stratified to ensure a uniform distribution of flow speeds and angles in both sets. A masked mean squared error (MSE) loss function was employed to account for variable trajectory lengths and padded waypoints as denoted by Eq. (19).
@@
where $y_{i}$ and $\hat{y}_{i}$ are the true and predicted waypoint coordinates, respectively, $m_{i}$ is the mask indicating valid waypoints ($m_{i}=1$) or padded positions ($m_{i}=0$), and $N$ is the total number of waypoint coordinates in the batch.
The network was trained using the Adam optimizer [31] with a learning rate of $1\times 10^{-4}$. Adam was chosen for its adaptive learning rate capabilities and robustness in training deep neural networks. The batch size was set to 64, and the model was trained for 50 epochs. Weight initialization was performed using the Xavier initialization method to ensure stable gradients at the start of training [32].
@@ 5.4 Regularization Techniques
To prevent overfitting, three regularization techniques were implemented:
  • Early Stopping: The validation loss was monitored, and training was halted if the loss did not improve for a set number of epochs (10).
  • Dropout: Although not included in the final model, dropout layers were experimented with during hyperparameter tuning but did not yield significant improvements.
  • L2 Regularization: L2 weight decay was applied implicitly via the Adam optimizer’s default settings.
@@ 5.6 Implementation Details
The neural network was implemented using PyTorch [33], a widely used deep learning framework that provides dynamic computation graphs and efficient GPU acceleration. Training was performed on a workstation equipped with an NVIDIA RTX2070 GPU to expedite the computational process.
@@
Compared to the traditional A* algorithms, the neural network offers significant computational speed-ups due to its feedforward nature, eliminating the need for iterative graph traversal. This efficiency is particularly advantageous for real-time applications where computational resources are limited and thus why it has been selected for analysis in this work.
@@ 5.7 Limitations and Considerations
5.8 Comparison with Related Work
Our approach aligns with recent trends in leveraging deep learning for path planning tasks [27, 28]. Unlike methods that use convolutional neural networks (CNNs) to process occupancy grids or reinforcement learning to learn navigation policies, our model directly predicts the sequence of waypoints, simplifying the integration into existing navigation systems.
@@ 6 Results
In this section, a comprehensive analysis of the performance of the four path planning approaches: the current-informed A* planner ($C.I._{A*}$), the current-informed neural network ($C.I._{NN}$), the wake-informed A* planner ($W.I._{A*}$), and the wake-informed neural network ($W.I.{NN}$) is provided. The evaluation metrics are the energy expenditure ($E$), path length ($L$), the number of high-velocity cells encountered ($n_{H,vel}$), turbulent cell count ($n_{turb.}$), and computational time ($t$) of each trajectory. The results are summarized in Table 1, with Fig. 4, 5 and 6 provided to further substantiate the results. The most optimal results are bolded for clarity in Table 1.
@@ 6.1 Energy Expenditure
Energy efficiency is critical in AUV operation due to their limited onboard energy resources. The wake-informed A* planner ($W.I._{A*}$) consistently demonstrates the lowest total energy expenditure across all speed ranges, as shown in Table 1. At the lowest speed range of 0.1–0.5m/s, $W.I._{A*}$ achieves an average energy expenditure of 51.49 joules with a standard deviation of 52.82 joules, outperforming the current-informed A* planner ($C.I._{A*}$) by approximately 11.3%. This persists at higher speeds, with $W.I._{A*}$ maintaining lower energy consumption compared to the other models. At the maximum tested speed range of 4.0–5.0m/s, $W.I._{A*}$ records an average energy expenditure of 11,952.30 joules, 7.2% less than $C.I._{A*}$.
@@ 6.2 Path Length
The path length metric ($L$) reflects the total distance traveled by the AUV along the planned trajectory. Shorter paths generally result in quicker LAR missions, for a given speed, but may traverse more hazardous areas. The current-informed A* planner ($C.I._{A*}$) achieves the shortest average path lengths across all speed ranges, with consistent trajectory lengths of 104.22–104.31m and a standard deviation of approximately 9.5m. This suggests that $C.I._{A*}$ is able to generate the most direct paths, but without considering wake structures and associated hazards. In contrast, the wake-informed A* planner ($W.I._{A*}$) shows slightly longer path lengths, averaging around 107 m, indicating that it deliberately avoids certain areas to minimize energy expenditure and exposure to high-velocity regions. The neural network models, particularly $C.I._{NN}$, exhibit significantly longer path lengths with higher variability. For example, $C.I._{NN}$ records an average path length of 133.27m in the 4.0–5.0m/s speed range, approximately 27% longer than $C.I._{A*}$. This suggests that the neural networks may not approximate the optimal path lengths as effectively, potentially leading to suboptimal routing. Fig. 5 provides an assessment of the median path length of the models against speed and angle. It can be seen that the A* current-informed method is the most consistent, with the A* wake-informed method also exhibiting similar performance. The neural networks result in a higher median path length, with suboptimal path generation at low speed and angles for the wake-informed NN.
@@ 6.3 High-Velocity Cells Encountered
Navigating through high-velocity cells poses increased risk and potential for energy expenditure, due to increased drag and control challenges associated with traversing the larger vehicle’s wake structure. The wake-informed A* planner ($W.I._{A*}$) exhibits the lowest average number of high-velocity cells encountered ($n{H,vel}$) across all speed ranges, as it has knowledge of the structures. At speeds between 0.5–1.0 m/s, $W.I._{A*}$ encounters an average of 2.73 high-velocity cells, significantly lower than $C.I._{A*}$ (7.32) on average. This indicates that $W.I._{A*}$ sufficiently avoids the high-velocity regions it is informed of, enhancing operational safety. Both neural network models exhibit less consistent performance in this metric. $C.I._{NN}$ is observed to reduce high-velocity cell encounters compared to $C.I._{A*}$, whilst $W.I._{NN}$ generally encounters more high-velocity cells than $W.I._{A*}$. This suggests that the neural networks may not fully capture the nuances of avoiding hazardous regions, possibly due to limitations in the training data or model capacity. Fig. 6 details the median number of high velocity cells encountered by each method. Similar performance is observed for the current-informed A* and NN, and the wake-informed NN. The wake-informed A* planner exhibits beneficial performance in comparison to the other models, with typically 0 - 5 cells encountered irrespective of speed and angle.
@@ 6.4 Turbulent Cell Count
The turbulent cell count ($n_{turb.}$) reflects the number of cells along the trajectory where any velocity fluctuation occurs, which would pose an additional control challenge for the AUV. It should be noted that no thresholding is applied to this, so small and large fluctuations are treated equally. The current-informed A* planner ($C.I._{A*}$) records the lowest turbulent cell counts across all speed ranges, with an average of 31.29 cells in the 0.5–1.0 m/s range. The wake-informed models, both $W.I._{A*}$ and $W.I._{NN}$, exhibit higher turbulent cell counts. This suggests that while the wake-informed planners effectively avoid high-velocity cells, they may traverse areas with higher turbulence. One possible explanation is that in avoiding high-velocity wake regions, the wake-informed planners route the AUV through areas with greater velocity disparities. The neural network models again display higher turbulent cell counts compared to their A* counterparts, indicating a potential area for improvement in modeling turbulence avoidance.
@@ 6.5 Computational Time
$n_{H,vel}$ $[cells]$ 0.1 - 0.5 2.27 $\pm$ 4.30 2.28 $\pm$ 3.45 0.83 $\pm$ 1.41 2.18 $\pm$ 3.71
@@
Table 1: Computational results for the planners and neural networks. Mean and standard deviation provided for the current-informed A* path planner ($C.I._{A*}$), current-informed neural network ($C.I._{NN}$), wake-informed A* path planner ($W.I._{A*}$) and wake-informed neural network ($W.I._{NN}$). Energy expenditure denoted by $E$, path length in meters by $L$, and the number of high velocity cells by $n_{H,vel}$ and the turbulent cell count by $n_{turb.}$. Time taken to compute a trajectory denoted by $t$.
Figure 6: Median count of the trajectories’ encountered high velocity cells against speed and inflow angle for the current-informed A* path planner ($C.I._{A*}$), current-informed neural network ($C.I._{NN}$), wake-informed A* path planner ($W.I._{A*}$) and wake-informed neural network ($W.I._{NN}$)
@@ 7 Discussion
The results demonstrate a trade-off between computational efficiency and path optimality. The wake-informed A* planner ($W.I._{A*}$) provides the most energy-efficient and safe paths but requires significant computational resources and time. Conversely, the neural network models offer near-instantaneous computation times, critical for real-time operations, however, the resulting paths are less optimal in terms of energy consumption and safety metrics, which may hinder their suitability. These findings highlight the importance of incorporating detailed environmental information, such as wake structures, into path planning algorithms. Although neural networks capture some aspects of the optimal paths, there is room for improvement in their ability to approximate the solutions provided by the A* planners.
@@
Previous studies have demonstrated the effectiveness of neural networks in path planning tasks for autonomous systems [27, 28]. However, most of these approaches focus on simplified environments or two-dimensional spaces. Our results extend this understanding by evaluating and demonstrating the potential of using neural network-based planners in complex three-dimensional underwater environments with dynamic wake effects.
Neural networks are able to reduce the computational time by 6 orders of magnitude, whilst resulting in paths that are only 4.51–19.79% less energy efficient and 9.81–24.38% longer, whilst encountering a similar, and sometimes lesser number of high velocity or turbulent cells on a given trajectory.
@@
The significant computational speed-ups offered by the neural networks, are well understood. This work shows that the extension to 3-dimensional planning aligns with expected findings in the current state of the literature [29]. The observed reductions in path optimality, however, also suggest that further advancements in neural network architectures or training methodologies could be beneficial to close the performance gap with traditional A* algorithms in complex, highly dynamic environments.

### v2

@@ 5.2 Model Architecture
Both neural network architectures are fully connected feedforward networks, designed to regress the sequence of waypoint coordinates directly from the input features. The architecture consists of an input layer, two hidden layers and an output layer as detailed in Fig. 5.
Figure 5: Neural network topology. Input to the network is the start location ($x_{s},y_{s},z_{s}$), goal location ($x_{g},y_{g},z_{g}$), flow speed ($u$) label and angle of attack label ($\theta$). The input vector has dimension (1, 8). This is then progressively increased by the linear layers to (1, 390), which is decomposed to a trajectory through ($[x_{1},y_{1},z_{1}],...,[x_{130},y_{130},z_{130}]$). This is then compared to the trajectories of the respective planner the NN is attempting to replicate ($W.I._{A*}$ or $C.I._{A*}$), before the loss is computed and backpropagated through the network.
Further details are provided in Table 7 in C. The choice of layer sizes was determined empirically to balance model capacity and computational efficiency. ReLU activation functions were used to introduce non-linearity and mitigate the vanishing gradient problem (Glorot et al., 2011). A fully connected architecture was chosen for its simplicity and to serve as a baseline, ensuring that future studies can easily compare advanced architectures (e.g., RNNs, Transformers) against this initial model.
@@ 5.3 Training Procedure
where $x$ is the feature value, $\mu$ is the mean, and $\sigma$ is the standard deviation computed from the training set. The dataset was then split into training and validation sets using an 80/20 split, resulting in 14,400 training samples and 3,600 validation samples. The split was stratified to ensure a uniform distribution of flow speeds and angles in both sets. A masked mean squared error (MSE) loss function was employed to score the ability of the two networks to predict wake-informed or current-informed trajectories. To determine the loss for a given path, the MSE loss was computed between the path generated by the neural network and the corresponding known path from the training set. This loss was then back-propagated through the network. The loss function is given by Eq. (17).
@@
where $y_{i}$ and $\hat{y}_{i}$ are the true and predicted waypoint coordinates, respectively, $m_{i}$ is the mask indicating valid waypoints ($m_{i}=1$) or padded positions ($m_{i}=0$), due to the variable trajectory length consisting of $N$ way-point coordinates in the path.
The network was trained using the Adam optimizer (Kingma and Ba, 2015) with a learning rate of $1\times 10^{-4}$. Adam was chosen for its adaptive learning rate capabilities and robustness in training deep neural networks. The batch size was set to 64, and the model was trained for 50 epochs. Weight initialization was performed using the Xavier initialization method to ensure stable gradients at the start of training (Glorot and Bengio, 2010).
@@ 5.4 Regularization Techniques
To prevent overfitting, three regularization techniques were implemented: (1) Early Stopping: The validation loss was monitored, and training was halted if the loss did not improve for a set number of epochs (10). (2) Dropout: Although not included in the final model, dropout layers were experimented with during hyperparameter tuning but did not yield significant improvements. (3) L2 Regularization: L2 weight decay was applied implicitly via the Adam optimizer’s default settings.
@@ 5.6 Implementation Details
The neural network was implemented using PyTorch (Paszke et al., 2019), a widely used deep learning framework that provides dynamic computation graphs and efficient GPU acceleration. Training was performed on a workstation equipped with an NVIDIA RTX2070 GPU to expedite the computational process. Further details, if required, are provided in Table 7 in C.
@@
Compared to the traditional A* algorithms, the neural network offers significant computational speed-ups due to its feedforward nature, eliminating the need for iterative graph traversal. Additionally, computation of the energy cost over the trajectory is not required - like in the A* planners - reducing the computational hardware requirement and energy consumption. This efficiency is particularly advantageous for real-time applications where computational resources are limited and thus why it has been selected for analysis in this work.
@@ 6 Results
This section presents a comparative performance analysis of the four path planning approaches developed and investigated in this study. These are:
  1. 1.
    Current-informed A* planner ($C.I._{A*}$)
  2. 2.
    Wake-informed A* planner ($W.I._{A*}$)
  3. 3.
    Current-informed neural network ($C.I._{NN}$)
  4. 4.
    Wake-informed neural network ($W.I.{NN}$)
The performance of each approach is evaluated based on several key metrics. These are the total energy expenditure of the generated paths ($E$) in Table 2, the total path length ($L$) in meters in Table 3, the number of high velocity ”risky” cells that the paths intersect ($N_{\text{high-velocity}}(P)$) in Table 4, the number of turbulent ”risky” cells that the paths intersect ($N_{\text{turbulent}}(P)$) in Table 4 and the time taken to generate a given path ($t$) in Table 5.
The aggregated results (mean and standard deviation across speed and angle combinations) are presented in Tables 2 - 5, where the optimal values have been highlighted. Figures 6, 7, and 8 provide further visual analysis of the results’ distribution and trends against flow speed and angle.
@@ 6.1 Energy Expenditure
Energy efficiency is critical in AUV operation due to their limited onboard energy resources. The wake-informed A* planner ($W.I._{A*}$) consistently demonstrates the lowest total energy expenditure, as determined via Eq. 14, across all speed ranges, as shown in Table 2. At the lowest speed range of 0.1–0.5m/s, $W.I._{A*}$ achieves an average energy expenditure of 51.49 joules with a standard deviation of 52.82 joules, outperforming the current-informed A* planner ($C.I._{A*}$) by approximately 11.3%. This persists at higher speeds, with $W.I._{A*}$ maintaining lower energy consumption compared to the other models. At the maximum tested speed range of 4.0–5.0m/s, $W.I._{A*}$ records an average energy expenditure of 11,952.30 joules, 7.2% less than $C.I._{A*}$.
@@ 6.2 Path Length
The path length metric ($L$), as given in Table 3, and determined via Eq. (15), reflects the total distance traveled by the AUV along the planned trajectory. Shorter paths generally result in quicker LAR missions, for a given speed, but may traverse more hazardous areas. The current-informed A* planner ($C.I._{A*}$) achieves the shortest average path lengths across all speed ranges, with consistent trajectory lengths of 104.22–104.31m and a standard deviation of approximately 9.5m. This suggests that $C.I._{A*}$ is able to generate the most direct paths, but without considering wake structures and associated hazards. In contrast, the wake-informed A* planner ($W.I._{A*}$) shows slightly longer path lengths, averaging around 107 m, indicating that it deliberately avoids certain areas to minimize energy expenditure and exposure to high-velocity regions. The neural network models, particularly $C.I._{NN}$, exhibit significantly longer path lengths with higher variability. For example, $C.I._{NN}$ records an average path length of 133.27m in the 4.0–5.0m/s speed range, approximately 27% longer than $C.I._{A*}$. This suggests that the neural networks may not approximate the optimal path lengths as effectively, potentially leading to suboptimal routing. Fig. 7 provides an assessment of the median path length of the models against speed and angle. It can be seen that the A* current-informed method is the most consistent, with the A* wake-informed method also exhibiting similar performance. The neural networks result in a higher median path length, with suboptimal path generation at low speed and angles for the wake-informed NN.
@@ 6.3 High-Velocity Cells Encountered
Navigating through high-velocity cells poses increased risk and potential for energy expenditure, due to increased drag and control challenges associated with traversing the larger vehicle’s wake structure. This data is provided in Table 4. The wake-informed A* planner ($W.I._{A*}$) exhibits the lowest average number of high-velocity cells encountered ($N_{\text{high-velocity}}(P)$) across all speed ranges, as it has knowledge of the structures. At speeds between 0.5–1.0 m/s, $W.I._{A*}$ encounters an average of 2.73 high-velocity cells, significantly lower than $C.I._{A*}$ (7.32) on average. This indicates that $W.I._{A*}$ sufficiently avoids the high-velocity regions it is informed of, enhancing operational safety. Both neural network models exhibit less consistent performance in this metric. $C.I._{NN}$ is observed to reduce high-velocity cell encounters compared to $C.I._{A*}$, whilst $W.I._{NN}$ generally encounters more high-velocity cells than $W.I._{A*}$. This suggests that the neural networks may not fully capture the nuances of avoiding hazardous regions, possibly due to limitations in the training data or model capacity. Fig. 8 details the median number of high velocity cells encountered by each method. Similar performance is observed for the current-informed A* and NN, and the wake-informed NN. The wake-informed A* planner exhibits beneficial performance in comparison to the other models, with typically 0 - 5 cells encountered irrespective of speed and angle.
@@ 6.4 Turbulent Cell Count
The turbulent cell count ($N_{\text{turbulent}}(P)$), as given in Table 4, reflects the number of cells along the trajectory where any velocity fluctuation occurs, which would pose an additional control challenge for the AUV. It should be noted that no thresholding is applied to this, so small and large fluctuations are treated equally. The current-informed A* planner ($C.I._{A*}$) records the lowest turbulent cell counts across all speed ranges, with an average of 31.29 cells in the 0.5–1.0 m/s range. The wake-informed models, both $W.I._{A*}$ and $W.I._{NN}$, exhibit higher turbulent cell counts. This suggests that while the wake-informed planners effectively avoid high-velocity cells, they may traverse areas with higher turbulence. One possible explanation is that in avoiding high-velocity wake regions, the wake-informed planners route the AUV through areas with greater velocity disparities. The neural network models again display higher turbulent cell counts compared to their A* counterparts, indicating a potential area for improvement in modeling turbulence avoidance.
@@ 6.5 Computational Time
Table 2: Energy expenditure (Eq. (14)) denoted by $E$ for the A* planners and neural networks. Mean and standard deviation provided for the current-informed A* path planner ($C.I._{A*}$), current-informed neural network ($C.I._{NN}$), wake-informed A* path planner ($W.I._{A*}$) and wake-informed neural network ($W.I._{NN}$).
Metric [$\downarrow$] Range [m/s] $C.I._{A*}$ $C.I._{NN}$ $W.I._{A*}$ $W.I._{NN}$
@@
Table 3: Path length (Eq. 15) in meters denoted by $L$ for the A* planners and neural networks. Mean and standard deviation provided for the current-informed A* path planner ($C.I._{A*}$), current-informed neural network ($C.I._{NN}$), wake-informed A* path planner ($W.I._{A*}$) and wake-informed neural network ($W.I._{NN}$).
Metric [$\downarrow$] Range [m/s] $C.I._{A*}$ $C.I._{NN}$ $W.I._{A*}$ $W.I._{NN}$
$N_{\text{high-velocity}}(P)$ 0.1 - 0.5 2.27 $\pm$ 4.30 2.28 $\pm$ 3.45 0.83 $\pm$ 1.41 2.18 $\pm$ 3.71
@@
Table 4: Number of high velocity cells $N_{\text{high-velocity}}(P)$ and the turbulent cell count $N_{\text{turbulent}}(P)$ (Eq. 10 & 12), encountered on a given trajectory $P$ by the A* planners and neural networks. Mean and standard deviation provided for the current-informed A* path planner ($C.I._{A*}$), current-informed neural network ($C.I._{NN}$), wake-informed A* path planner ($W.I._{A*}$) and wake-informed neural network ($W.I._{NN}$).
Figure 8: Median count of the trajectories’ encountered high velocity cells ($N_{\text{high-velocity}}(P)$) against speed and inflow angle for the current-informed A* path planner ($C.I._{A*}$), current-informed neural network ($C.I._{NN}$), wake-informed A* path planner ($W.I._{A*}$) and wake-informed neural network ($W.I._{NN}$)
Metric [$\downarrow$] Range [m/s] $C.I._{A*}$ $C.I._{NN}$ $W.I._{A*}$ $W.I._{NN}$
@@
Table 5: Computational time for a given path by the A* planners and neural networks. Mean and standard deviation provided for the current-informed A* path planner ($C.I._{A*}$), current-informed neural network ($C.I._{NN}$), wake-informed A* path planner ($W.I._{A*}$) and wake-informed neural network ($W.I._{NN}$).
@@ 7 Discussion
The results demonstrate a trade-off between computational efficiency and path optimality. Knowledge of the wake, via the wake-informed A* planner ($W.I._{A*}$), provides the most energy-efficient and safe paths but requires significant computational resources and time. Conversely, the neural network models offer near-instantaneous computation times, critical for real-time operations, however, the resulting paths are less optimal in terms of energy consumption and safety metrics, which may hinder their suitability. These findings highlight the importance of incorporating detailed environmental information, such as wake structures, into path planning algorithms. Although neural networks capture some aspects of the optimal paths, there is room for improvement in their ability to approximate the solutions provided by the A* planners.
@@
Previous studies have demonstrated the effectiveness of neural networks in path planning tasks for autonomous systems (Tai et al., 2017; Pfeiffer et al., 2017). However, most of these approaches focus on simplified environments or two-dimensional spaces. Our results extend this understanding by evaluating and demonstrating the potential of using neural network-based planners in complex three-dimensional underwater environments with dynamic wake effects.
The observed reduction in path optimality (higher energy cost, longer paths) when using the NN approximations carries significant implications. For missions operating under tight energy budgets or requiring precise maneuvering near hazards, the 4.51–19.79% increase in energy expenditure or deviations from the truly optimal path might be unacceptable. This highlights a critical challenge: while NNs offer the computational speed needed for real-time reactivity, their current level of precision might limit their applicability in safety- or efficiency-critical scenarios without further safeguards or improvements. The approximation errors likely arise from the inherent difficulty for a feedforward network to perfectly replicate the complex, iterative decision-making process of A* within a high-dimensional, dynamic 3D flow field, especially given the potentially non-linear relationships between inputs and optimal outputs.
While the dramatic reduction in computation time is the most prominent advantage of the NN approach demonstrated here, other potential benefits, do exist. NNs offer a fixed, predictable inference time once trained, unlike A* whose runtime can vary considerably depending on the specific start/goal configuration and environment complexity. If trained on a sufficiently diverse dataset encompassing various environmental conditions and vehicle states, NNs might exhibit smoother interpolation and potentially better generalization to unforeseen (but similar) scenarios compared to A*’s discrete grid search. It should be noted that this adaption would requires rigorous validation against out-of-distribution data, potentially from a much larger set of paths $P$. Furthermore, the NN framework is inherently adaptable; it could potentially incorporate diverse real-time sensor inputs (e.g., sonar, Doppler Velocity Log) beyond the pre-defined parameters used here, allowing for more reactive planning based on immediate perception – and potentially tailored to specific vehicle configurations. Although not explicitly optimized for here, certain NN architectures could potentially learn to generate inherently smoother paths, which might be more dynamically feasible for an AUV to follow, reducing control effort compared to potentially jagged A* paths on a grid. However, based on the results here, comparing the optimality metrics, it is the computational speed, compute hardware and energy requirement reduction that remains the primary realized advantages of the NN approach in this work.
Deploying either the A* or NN path planners in real-world AUV operations involves significant practical complications beyond the idealized simulation environment. For A*, the primary hurdles are its computational demand which challenges real-time re-planning on typical AUV onboard processors, especially in dynamic environments, and its reliance on a pre-defined, accurate environmental map. Real-world sensor data (e.g., sonar for obstacle detection, ADCP for currents) is often noisy, sparse, and provides incomplete coverage, making the creation of a reliable, up-to-date grid map difficult. Furthermore, the discretization of the continuous world into a grid can itself introduce sub-optimality or miss feasible paths.
For NNs, despite their online speed, the main concerns are reliability and safety. Their performance is fundamentally tied to the quality and representativeness of the training data. Encountering real-world conditions significantly different from the training scenarios (out-of-distribution problem) can lead to unpredictable and potentially unsafe behavior. Verification and validation of NN controllers for safety-critical applications like AUV navigation remain open research challenges. Both methods, as implemented here, also simplify AUV dynamics, neglecting constraints like minimum turning radius, acceleration/deceleration limits, and control response times, which are crucial for path feasibility. Additionally, handling unforeseen dynamic obstacles (e.g., marine life, other vessels) requires integration with reactive collision avoidance systems, which is not covered in the current planning framework. Bridging the gap between these simulation-based planners and robust real-world deployment requires addressing sensor integration, uncertainty management, dynamic replanning, vehicle constraints, and rigorous safety validation.
Several avenues exist for potentially improving the precision and reliability of NN-based planners in this context. Firstly, exploring more sophisticated network architectures tailored to sequential decision-making, such as Recurrent Neural Networks (RNNs, e.g., LSTMs) or Transformers (Vaswani et al., 2017), could better capture the temporal dependencies along the path and potentially model the environmental interactions more effectively. Secondly, integrating physics knowledge directly into the network, perhaps through Physics-Informed Neural Networks (PINNs) (Raissi et al., 2019) or by designing loss functions that explicitly penalize physically implausible or high-energy path segments, might constrain the network to produce more realistic and efficient trajectories. Thirdly, enhancing the training strategy, for instance by using Generative Adversarial Networks (GANs) to generate paths or employing reinforcement learning (RL) where the agent learns a policy directly by interacting with the environment or a simulator, could lead to better performance, although RL often requires substantial training time and careful reward shaping. Finally, hybrid approaches, where an NN provides a fast initial path suggestion that is then locally refined using a quicker optimization or search algorithm (like a limited-depth A* or potential field method), could offer a pragmatic balance between speed and near-optimality.
@@
The significant computational speed-ups offered by the neural networks, are well understood. This work shows that the extension to 3-dimensional planning aligns with expected findings in the current state of the literature (LeCun et al., 2015). The observed reductions in path optimality, however, also suggest that further advancements in neural network architectures or training methodologies could be beneficial to close the performance gap with traditional A* algorithms in complex, highly dynamic environments.